# Birack Bracket Quivers and Framed Links

### v1

@@ 2 Birack Review
As mentioned in the Introduction, to define knot invariants various mathematical structures have been implemented to encode the crossing information of a knot, taking the forms of groups, tensors, and more abstract algebraic structures. Reidemeister moves characterize the equivalence relation of ambient isotopy on diagrams; the corresponding algebraic structures in result in different knot invariants. More specifically, quandles are represent the algebraic structure of Reidemeister moves on arcs in classical knots; biquandles represent Reidemeister moves on semi-arcs in classical and virtual knots, and racks represent the algebraic structure of arcs in framed knots and links through replacing Reidemeister I move with the framed version of the move. In this paper, we consider the case of biracks, algebraic structures generated by semiarcs in a link diagram with axioms corresponding to blackboard framed isotopy.
Recall that two blackboard framed oriented knot or link diagram represent isotopic framed knots or links if and only if they are related by a sequence of the blackboard framed Reidemeister moves, of which a generating set are:
@@
    The maps $\alpha_{x},\beta_{x}:X\to X$ and $S:X\times X\to X\times X$ defined by $\alpha_{x}(y)=y\,\overline{\triangleright}\,x$, $\beta_{y}(x)=y\,\underline{\triangleright}\,x$ and $S(x,y)=(y\,\overline{\triangleright}\,x,x\,\underline{\triangleright}\,y)$ are invertible for all $x\in X$ and
@@
Biracks are sometimes defined without requiring our axiom (i); we will refer to these as regular biracks. Essentially, regular biracks can be used to define invariants of framed isotopy, e.g. braid invariants, while framed biracks provide invariants of framed knots and links. Unless otherwise specified, we will mean “framed birack” whenever we say “birack”.
@@
The birack axioms are chosen so that for every birack coloring of a framed oriented link diagram before a framed Reidemeister move, there is a unique coloring of the diagram after the move that agrees with the original coloring outside the neighborhood of the move.
@@ 3 Birack Brackets
Let $X$ be a birack and $R$ a commutative unital ring. A birack bracket is defined by two maps $A,B:X\times X\to R^{\times}$ denoted $A(x,y)=A_{x,y}$ and $B(x,y)=B_{x,y}$ such that
@@ 4 Birack Bracket Quivers
As in previous cases such as [3, 5, 7] etc., we observe that since the ingredients for $\mathcal{Q}_{X,S}(L)$ are all invariant under framed Reidemeister moves, so is $\mathcal{Q}_{X,s}(L)$. That is, we have:
@@
For any finite biquandle $X$ and $S\subseteq\mathrm{Hom}(X,X)$, the quiver $\mathcal{Q}_{X,S}(L)$ is an invariant of framed isotopy.
@@
Given a finite birack $X$, a framed knot $K$ and a birack bracket $\beta$ with coefficients in a commutative unital ring $R$ and $S\subset\mathrm{Hom}(X,X)$ a set of endomorphisms of $X$. Then the birack bracket quiver $\mathcal{BRQ}_{X,S}^{\beta}(K)$ has vertex set $V$ and edge set $E$. Let $P$ the set of maximal non-repeating paths in the quiver. We then decategorify this birack bracket quiver $\mathcal{BRQ}_{X,S}^{\beta}(K)$ to obtain the birack bracket in-degree polynomial

### v2

@@ 2 Birack Review
As mentioned in the Introduction, to define knot invariants various mathematical structures have been implemented to encode the crossing information of a knot, taking the forms of groups, tensors, and more abstract algebraic structures. Reidemeister moves characterize the equivalence relation of ambient isotopy on diagrams; the corresponding algebraic structures in result in different knot invariants. More specifically, quandles represent the algebraic structure of Reidemeister moves on arcs in classical knots; biquandles represent Reidemeister moves on semi-arcs in classical and virtual knots, and racks represent the algebraic structure of arcs in framed knots and links through replacing Reidemeister I move with the framed version of the move. In this paper, we consider the case of biracks, algebraic structures generated by semiarcs in a link diagram with axioms corresponding to blackboard framed isotopy.
Recall that two blackboard framed oriented knot or link diagram represent isotopic framed knots or links if and only if they are related by a sequence of the blackboard framed Reidemeister moves, of which a generating set is:
@@
    The maps $\alpha_{x},\beta_{x}:X\to X$ and $S:X\times X\to X\times X$ defined by $\alpha_{x}(y)=y\,\overline{\triangleright}\,x$, $\beta_{y}(x)=x\,\underline{\triangleright}\,y$ and $S(x,y)=(y\,\overline{\triangleright}\,x,x\,\underline{\triangleright}\,y)$ are invertible for all $x\in X$ and
@@
Biracks are sometimes defined without requiring our axiom (i); we will refer to these as regular biracks. Essentially, regular biracks can be used to define invariants of regular isotopy, e.g. braid invariants, while framed biracks provide invariants of framed knots and links. Unless otherwise specified, we will mean “framed birack” whenever we say “birack”.
@@
More precisely, the framed birack axioms are chosen so that for any valid birack coloring of an oriented framed link diagram on one side of a framed Reidemeister move, there is a unique valid coloring of the diagram on the other side of the move which agrees outside the neighborhood of the move. Then starting with a single strand, using Reidemeister II and III moves we can perform what’s sometimes called the Fenn-Rourke move (FR):
[…]
The bijectivity of colorings establishes the bijectivity of the kink map $\pi$; the framed Reidemeister I move condition expressed algebraically by axiom (i) then shows that the kink map (and hence its inverse) does not depend on which side of the strand the loop is on.
@@ 3 Birack Brackets
Let $X$ be a birack and $R$ a commutative unital ring. A birack bracket is defined by two maps $A,B:X\times X\to R^{\times}$ from ordered pairs of elements of $X$ to units in $R$ denoted $A(x,y)=A_{x,y}$ and $B(x,y)=B_{x,y}$ such that
@@ 4 Birack Bracket Quivers
As in previous cases such as [3, 5, 7] etc., we observe that since the ingredients for $\mathcal{BRQ}_{X,S}(L)$ are all invariant under framed Reidemeister moves, so is $\mathcal{BRQ}_{X,s}(L)$. That is, we have:
@@
For any finite birack $X$ and $S\subseteq\mathrm{Hom}(X,X)$, the quiver $\mathcal{Q}_{X,S}(L)$ is an invariant of framed isotopy.
@@
Given a finite birack $X$, a framed knot $K$ and a birack bracket $\beta$ with coefficients in a commutative unital ring $R$ and $S\subset\mathrm{Hom}(X,X)$ a set of endomorphisms of $X$. Then the birack bracket quiver $\mathcal{BRQ}_{X,S}^{\beta}(K)$ has vertex set $V$ and edge set $E$. Let $P$ the set of maximal non-repeating paths (i.e., ordered sequences of edges with each edge terminating at the initial vertex of the next edge such that no edge is included more than once and such that the path is not a sub-path of any other path) in the quiver. We then decategorify this birack bracket quiver $\mathcal{BRQ}_{X,S}^{\beta}(K)$ to obtain the birack bracket in-degree polynomial